Simultaneous equations
X+y=2
X(x-y)=12

Answer by oscargut(2103)

X+y=2
X(x-y)=12

y = 2-x
x(x-(2-x))=12
x(2x-2)=12
2x(x-1)=12
Answer: x = 3 and y = -1 or x = -2 and y = 4
I can help you with more questions at: mthman@gmail.com
Thanks